A motorcycle begins driving from at rest position. It accelerates at 8 km/hr2 for 20 seconds, then travels at a steady speed for another 20 seconds, all in the same direction. How much distance has it covered since it started driving?



1
Expert's answer
2022-06-02T15:44:25-0400

"s=\frac{at^2}2+at\cdot t=1.5at^2=18.5~mm."